OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. 8,3) m=6 can someone help me please

OpenStudy (anonymous):

Point slope form is Y- Y1= M(X-X1)

OpenStudy (anonymous):

So to fill it in it would be Y-3=6(x-8)

OpenStudy (anonymous):

oh okay thanks Anecia23 that really helped

OpenStudy (anonymous):

No problem! :)